# Verb Mirage: Unveiling and Assessing Verb Concept Hallucinations in Multimodal Large Language Models

### v1

@@ 4.1 Token Uncertainty
Finding: MLLM verb hallucination has something in common with MLLM object hallucination. Uncertainty is strongly related to MLLM hallucination.
Figure 7: Probability distribution of hallucinated tokens and not hallucinated tokens.
4.2 Attention
4.2.1 Key Image Area Attention
We hypothesize that models pay less attention to key information in images and text when they give hallucinated answers. We still take LLaVA V1.5 as an example and study the relationship between LLaVA V1.5 verb hallucination and its attention to tokens related to key areas in images. From Fig. 8(a), we can see that there is an obvious distinction of distribution between hallucinated attention and non-hallucinated attention, showing a strong correlation between inadequate attention to key areas and hallucination.
4.2.2 Verb Token Attention
Previous experiments give us a strong impression that MLLM relies too much on objects. Therefore, we hypothesize that MLLM tends to pay more attention to object tokens. In Fig. 8(b), MLLM’s attention to verb tokens is normalized by the sum of verb attention and object attention. As MLLM tends to answer Yes, we focus on questions with correct answers No. For negative YN questions with object references, hallucinated tokens are given with inadequate attention to the verb in question. In Fig. 8(c), verb and noun attention are normalized by their sum. MLLM tends to give more attention to object tokens than verb tokens.
Finding: MLLM tends to pay more attention to object tokens than verb tokens.
Figure 8: Attention for YN questions with object references.
4.3 MLLM Error Consistency
Consistency on MC and YN questions. From Tab. 1, we find that there is inconsistency in MLLMs’ performance on multiple-choice and yes-or-no tasks. However, it is an interesting question whether there is some consistency. Thus, we divide the questions into different classes according to object class references and get the object classes that influence MLLMs the most. Assume $S_{1}$ denotes the set of object classes without which MLLMs can not recall correct verbs given YN questions, and $S_{2}$ denotes the similar set of object classes for MC questions. We define overlap index=$\frac{|S_{1}\cap S_{2}|}{\text{min}\{|S_{1}|,|S_{2}|\}}$, and compute MLLMs’ overlap index. From Tab. 4, we find substantial overlap between $S_{1}$ and $S_{2}$. A substantial number of object classes can affect MLLM verb understanding of both question types.
Qwen-VL-Chat mPLUG-Owl2 LLaVA V1.5 InstructBLIP
overlap 95.35 79.17 82.14 100
Table 4: Influential object set overlap of MLLMs.
Consistency among models. We analyze models and get verb classes that are influenced by object references most according to per verb class Yes ratio. From Tab. 5, we can see that there is substantial overlap among different models. We also ensemble the answers given by models and show results in Fig. 1. There is no substantial improvement over the individual models, showing that the models share similar biases.
Finding: A substantial number of objects affect MLLM verb understanding on both question formats and all models. MLLMs show error consistency to some extent.
IoU Qwen-VL-Chat LLaVA V1.5 mPLUG-Owl2 InstructBLIP
Qwen-VL-Chat - 59.30 42.86 52.78
LLaVA V1.5 59.30 - 46.91 53.57
mPLUG-Owl2 42.86 46.91 - 41.54
InstructBLIP 52.78 53.57 41.54 -
Table 5: YN influenced verb set IOU.
5 Hallucination Mitigation Methods
5.1 Training-Free Methods
@@
OPERA [huang2024opera], VCD [leng2024mitigating], and HALC [chen2424halc] are outstanding hallucination mitigation methods working during inference. They do not require finetuning, so they are low-cost and have more general applicability. We evaluate them on benchmarks formed by MC and YN questions.
We present the results of OPERA in Tab. 6, showing that OPERA does not always show improvement. Even when it shows improvement, the gain is marginal. Though it has been claimed that OPERA tries to punish overreliance on summary tokens and force more attention on visual tokens [huang2024opera], from Fig. 9 we find that for questions with correct answer No hallucinated models tend to give more attention to visual tokens. If the reward is given for attention to visual tokens, hallucination will be worsened. As we can not know the correct answer ahead of time, we can not decide whether to reward high attention to visual tokens.
Figure 9: Probability distribution of visual token attention for positive samples and negative samples.
@@
VCD regards language prior as a hallucination-inducing factor, uses visual distortion to trigger MLLM’s reliance on language prior, and proposes contrastive decoding to eliminate hallucination. We present the results of VCD in Tab. 6. VCD shows an inconsistent effect on three models and thus fails on our benchmarks. To dig deeper, we compute Qwen-VL’s KL divergence between the original vocabulary distribution and contrasted vocabulary distribution $\text{KL}(p_{\theta}(y_{t}|v,x,y_{<t})||p_{vcd}(y_{t}|v,v^{\prime},x,y_{<t})))$ of 20, 000 samples, and find that the KL divergence of 18, 624 samples is 0, meaning that in our benchmarks there is no strong element of language prior which usually hides but can be triggered by visual distortion.
HALC argues that field of view is a strong factor affecting MLLM hallucination, and proposes contrastive decoding given different fields of view. As HALC is very time-consuming, we evaluate part of our benchmarks and present the results of HALC in Tab.  7. It also fails to show substantial improvement. On one hand, HALC focuses too much on objects. In our benchmarks, to minimize relevance to object hallucination, all object references are correct which limits the use of HALC. On the other hand, HALC’s failure shows that fields of view do not have much influence on MLLM verb understanding.
w/ obj w/o obj
YN acc vanilla 78.06 49.5 72.53 79.24 58.21 73.82
w/ OPERA 78.03 42.46 73.08 79.19 54.45 74.01
w/ VCD 77.23 52.38 73.88 78.88 57.86 74.39
YN prec vanilla 62.37 39.8 55.79 65.09 44.49 57.25
w/ OPERA 62.32 36.76 56.38 64.98 42.38 57.45
w/ VCD 61.42 41.18 57.42 65.48 44.27 58.28
YN recall vanilla 87.02 98.94 86.77 82.68 97.49 87.79
w/ OPERA 87.07 99.54 86.85 82.81 98.33 87.62
w/ VCD 86.11 98.3 85.3 78.83 97.26 84.28
MC acc vanilla 55.95 57.37 13.48 54.57 51 6.25
w/ OPERA 55.63 57.28 13.96 54.17 51.13 6.28
w/ VCD 55.95 54.26 5.77 52.47 48.94 3.74
Table 6: Comparison on benchmarks w/ and w/o OPERA and VCD.
YoN verb only (20K samples) MC verb only (20K samples)
w/o HALC w/ HALC w/o HALC w/ HALC
acc prec recall acc prec recall acc acc
mPLUG-Owl-2 55.85 42.47 97.74 56.07 42.59 97.76 60.18 60.18
LLaVA V1.5 54.04 41.58 99.25 54.00 41.56 99.25 51.26 51.26
InstructBLIP 72.44 55.05 86.78 72.44 55.05 86.78 6.54 6.54
YoN w/ obj (20K samples) MC w/ obj (20K samples)
w/o HALC w/ HALC w/o HALC w/ HALC
acc prec recall acc prec recall acc acc
mPLUG-Owl2 62.53 46.54 95.99 62.64 46.61 95.86 61.78 61.78
LLaVA V1.5 52.04 40.53 99.05 52.03 40.53 99.05 56.92 56.92
InstructBLIP 73.98 57.12 82.88 73.98 57.12 82.88 18.26 18.26
Table 7: Comparison on benchmarks w/ or w/o HALC.
5.2 Mitigation via Finetuning
The above experiments reveal severe MLLM verb hallucination and show very different properties from noun hallucination: there is no strong language prior, no strong reliance on visual tokens, and no strong dependence on fields of view. These phenomena give us a strong impression that MLLM did not learn verbs well. To mitigate verb hallucination to some extent without sacrificing MLLM’s ability in other perspectives, we further explore verb hallucination mitigation via direct finetuning and propose a method based on verb structure knowledge.
@@ 5.2 Mitigation via Finetuning
We try to advance the MLLM with Pangea [li2024isolated], which organizes existing heterogeneous action datasets in a unified way. It builds a mapping from action labels to abstract verb semantics. 290 frequent verb nodes in VerbNet [verbnet] are selected and a one-to-290 mapping is built. It gives us a whole picture of diverse verbs and carries the structure knowledge of verb relationships. We select 60,000 samples from Pangea according to the proportion of the source dataset and build an instruction-tuning dataset. It contains 280 out of 290 nodes in Pangea P2S mapping and covers a wide range of verb semantics. To improve MLLM’s ability on verb understanding without sacrificing MLLM’s ability on other perspectives, we choose to finetune LLaVA V1.5 with LoRA [hu2022lora]. The results are shown in Tab. 8.
From the result, we can see outstanding improvement in MC questions. Although Pangea only contains rough action labels, it proves helpful to verb hallucination mitigation. In the future, mining more action data according to the structured verb semantics and finetuning MLLMs on them can be a promising way to mitigate verb hallucination.
Finding: Finetuning model with data roughly annotated but containing rich verb semantics proves helpful to MLLM verb hallucination.
not Finetuned Finetuned not Finetuned Finetuned
YN YN MC MC
acc prec recall acc prec recall acc acc
w/ obj 54.75 42.37 97.31 48.24 39.21 98.99 52.60 62.20
verb only 75.49 59.64 84.14 61.00 46.23 97.11 50.42 62.17
Table 8: Finetuning result of LLaVA V1.5.
6 Discussion
We mainly probe verb hallucination via MC and YN questions and leave free-form captioning out, because our experiments show that MLLMs tend to give as few verbs as possible. Thus, evaluating verb hallucination via metrics similar to CHAIR score [rohrbach2018object] is impractical: unmentioned existing objects will not cause punishment. Only mentioned non-existing objects cause punishment. Metrics like the CHAIR score can not provide a full perspective for verb hallucination. Another issue is the synonym of verb/action and hierarchy processing. The same action can be described in many different ways(e.g., doing housework and mopping floor), and establishing such synsets is still a problem. Based on these considerations, we do not choose free-form captioning as verb hallucination probation.
Given the above experiments, verb hallucination is quite different from object hallucination. Although MLLMs have shown some understanding of verbs, their performance is still far from satisfactory. Existing training-free hallucination mitigation methods show inconsistent results and can be said to fail on our benchmarks. It is still an open question what factor inside MLLMs can trigger MLLM verb hallucination. In our experiments, verb understanding does not always rely on visual tokens, and there is no strong language prior. Given these observations, we hypothesize that existing widely used MLLMs overfit on objects but underfit on verbs. In the future, study about MLLM visual representation can give us concrete proof.
Mitigating verb hallucination via finetuning is a more promising way. As the data collection is labor-intensive and finetuning consumes high resources, how to fully use existing data is vital but still a problem to be explored. Pangea contains rich verb semantics and provides a guideline for data collection and selection. In the future, a full use of Pangea P2S mapping can provide guidance. As existing datasets also contain annotations about verbs, maybe we can delicately reuse data so that we can mitigate verb hallucination without introducing a large amount of data.
7 Conclusion
@@
In this paper, to the best of our knowledge, we first reveal MLLM verb hallucination and build a benchmark to probe it from various perspectives. Our experiment reveals that MLLMs suffer from severe verb hallucination in many ways, and existing training-free hallucination mitigation methods fail. The experiment shows that MLLM verb hallucination is quite different from object hallucination. Thus, we propose to finetune MLLM with LoRA using data of low quality but with rich verb semantics to mitigate verb hallucination. However, how to finetune existing models efficiently is still a problem to be explored. Moreover, whether there are more effective training-free verb hallucination mitigation methods is still an open problem.
@@ Appendix A Details on Benchmark Construction
A.1 Benchmark Based on HICO
HICO [chao2015hico] test set contains 9.8K images. We use the HICO test set for benchmark construction. It includes 600 action classes formed by 80 object classes and 117 verb classes. There is a special verb class “no_interaction”, and we leave out annotations about it.
@@ A.1 Benchmark Based on HICO
For MC question construction, given an image $I$ and its HOI annotation $\mathcal{H}_{I}$, we select $h_{0}=(v_{0},o_{0})\in\mathcal{H}_{I}$ as a positive option, select 3 negative options from the set of plausible but not presented action classes $\{(v,o)|o=o_{0}\land(v,o)\notin\mathcal{H}_{I}\}$, and form a MC question with one correct option. In our benchmark, there are 16.6K unique MC questions.
@@
A.2 Benchmark Based on CharadesEgo
CharadesEgo [sigurdsson2018actor] contains 157 action classes, each of which is formed by a verb and an object. However, many object classes in CharadesEgo can not match more than 2 verbs. Therefore, to build negative options, we sample from the set $\{(v,o)|o\in\mathcal{O}_{I}\land(v,o)\notin\mathcal{H}_{I}\}$, where $\mathcal{O}_{I}=\{o|(v,o)\in\mathcal{H}_{I}\}$. The ratio of questions with correct answers Yes and No is approximately 1:1. There are 76K unique YN questions and 36K unique MC questions, and we use 28K YN questions and 10K MC questions for experiment in the main paper.
@@ B.1 Uncertainty
of LLaVA V1.5 [liu2024improved]. In our setting, $y_{0}$’s are usually option tokens or Yes/No tokens, which are used to judge the correctness of MLLMs’ answers. Here we provide some additional visualizations in Fig. 10. The visualizations give us more solid proof that uncertainty is strongly related to MLLM verb hallucination.
Figure 10: Token uncertainty of InstructBLIP [dai2023instructblip]. Not hallucinated tokens are often given with high probability.
B.2 Key Image Area Attention
LLaVA V1.5 adopts CLIP [radford2021learning] as the visual encoder. 576 tokens generated by CLIP are used as visual tokens and fed to the language model after projection. Each of these 576 tokens represents a small area of the original image. Using the bounding box annotation of HICO, we can get the key image area of a certain verb, represented by the area bounded by the bounding boxes of humans and objects. Our visualization shows that hallucinated MLLM pays less attention to the key image area, but the margin is not large enough. Although MLLM can pay attention to the correct key areas, it can not often understand verb concepts.
B.3 Verb Token Attention
@@
B.4 Visual Token Attention
Visual Token Attention(VTA) is defined similarly to Visual Modality Contribution(VMC) [chen2024multi]. The Visual Token Attention is defined as
@@ B.4 Visual Token Attention
where $\alpha_{ij}$ represents the attention weight assigned to token $i$ at head $j$ in the last transformer layer, $V$ denotes the set of visual tokens, and $T$ denotes the set of textual tokens. In the implementation of OPERA [huang2024opera], when generating the first 10 tokens, a reward is given for high visual token attention. We visualize VTA on other models and other question types in Fig. 11 and  12, and have the following observations:
@@
B.5 Object Reliance
We split YN question sets according to object classes, and show the difference between mPLUG-Owl2’s [ye2024mplug] accuracy with and without object references in Fig. 13. We can see that for questions about most object classes, mPLUG-Owl2 behaves differently with and without object references and thus is affected by object references although the accuracy does not change much (62.94/62.61).

### v3

@@ Introduction
In conclusion, MLLM verb hallucination has something in common with MLLM object hallucination. Uncertainty is strongly related to MLLM verb hallucination.
Figure 9: Distribution of token uncertainty.
mAP vs. Acc
A natural question is, does low accuracy mean poor verb understanding? For example, in Tab. 1, LLaVA V1.5 shows an accuracy of 52.16 on YN questions with object references. To answer this question, we compare its performance with a strong baseline: HICO-finetuned CLIP (radford2021learning) with outstanding results (li2024isolated). To rule out the influence of objects, we test mAP under the “Known Object” setting (chao2015hico). We use the probability of Yes tokens to calculate mAP. LLaVA V1.5 achieves an mAP of 68.41 while HICO-finetuned CLIP has an mAP of 60.45. The detailed AP gap of each HOI class is in Fig. 11. LLaVA V1.5 outperforms CLIP in most HOI classes. This means that LLaVA V1.5, although showing severe verb hallucination, does know well which verbs an image is more likely to contain. Therefore, one of the sources of verb hallucination is the mis-calibration of tokens.
Figure 10: The influence of crop size on token uncertainty.
Field of View and Uncertainty
Previous research (chen2424halc; zhang2025mllms) has revealed the use of field of view ensembling in hallucination mitigation. (zhang2025mllms) has shown the influence on the change of token uncertainty. We also try to analyze the effect of field of view cropping on MLLM verb understanding. We select from HICO a set of images. For each image, there is only one HOI instance (to rule out the disturbance of multiple instances). The HOI instance is also small enough so that cropped images are substantially different from the original images. The token uncertainty on MC questions visualized in Fig. 10 shows that the uncertainty and accuracy do not change much: different from object, verb understanding is not substantially affected by the field of view: MLLMs do not always know the verb concepts in the area.
Hallucination Mitigation Methods
Figure 11: AP Comparison of LLaVA V1.5 and CLIP.
Training-Free Methods
@@
OPERA (huang2024opera), VCD (leng2024mitigating), and Nullu (yang2025nullu) are outstanding training-free hallucination mitigation methods. They do not require fine-tuning and thus are low-cost and have more general applicability.
We present the results of OPERA in Tab. 5, showing the marginal or even negative effect of OPERA. Though OPERA tries to punish overreliance on summary tokens and force more attention on visual tokens, we have discussed in Sec. Vision-Language Interaction that more attention to visual tokens does not exclude hallucination. If the reward is given for attention to visual tokens, hallucination may be worsened.
@@
VCD regards language prior as a hallucination-inducing factor, uses visual distortion to trigger it, and proposes contrastive decoding to mitigate hallucination. The results of VCD are in Tab. 5. VCD shows inconsistent effects on three models and thus fails on our benchmarks. To dig deeper, we compute Qwen-VL’s KL divergence between the original token distribution $p_{\theta}(y_{t}|v,x,y_{<t})$ and contrasted token distribution $p_{vcd}(y_{t}|v,v^{\prime},x,y_{<t}))$ of 20K samples, and find that the KL divergence of 18.6K samples is 0, meaning that MLLMs rely heavily on language prior to understand verbs, and such prior can not be easily omitted.
Nullu identifies a low-rank hallucination subspace from truthful and hallucinated responses, projects model weights into its null space to suppress false priors, and reduces hallucination without high cost. Its negative results in Tab. 5 show that MLLM layers have not formed reliable distinguishment between truthful and hallucinated verbs.
Influence of Fine-tuning
@@
REVERIE (zhang2024reflective), Haloquest (wang2024haloquest), and Octopus (suo2025octopus) are outstanding fine-tuning methods, and show decent results on object-centric benchmarks. Octopus train models to adopt different contrastive decoding methods in different cases, while the other two methods fine-tune models with meticulously designed training sets. However, these training sets do not contain much verb knowledge, and we show their limitation in Tab. 5. To mitigate verb hallucination to some extent without sacrificing MLLM’s ability in other perspectives, we explore the effect of datasets with rich verb knowledge. We try to advance the MLLM with Pangea (li2024isolated), which organizes existing heterogeneous action datasets in a unified way. It builds a mapping from action labels to abstract verb semantics. 290 frequent verb nodes in VerbNet (verbnet) are selected and a one-to-290 mapping is built. It gives us a whole picture of diverse verbs and carries the structure knowledge of verb relationships. We select 60K samples from Pangea according to the proportion of the source dataset and build an instruction-tuning dataset. Details are in Suppl. Sec. 4. It contains 280 out of 290 nodes in Pangea P2S mapping and covers a wide range of verb semantics. Following common practice (zhang2024reflective; wang2024haloquest), we fine-tune LLaVA V1.5 with LoRA. The results are shown in Tab. 5. Although Pangea only contains rough action labels, it proves more helpful to verb hallucination mitigation than previous training sets. The performance of the original/fine-tuned model on MMMU (yue2024mmmu): 32.2/33.0, MathVista (lu2023mathvista): 23.6/24.5. The result is not seriously impaired. Our method may also be integrated into MoE in case of interference. In the future, mining more action data according to the structured verb semantics and fine-tuning MLLMs on them can be a promising way to mitigate verb hallucination.
YN w/ obj YN w/o obj MC w/ obj MC w/o obj
acc F1 acc F1 acc acc
LLaVA V1.5 52.16 57.69 59.16 61.79 57.37 51.00
OPERA 42.46 53.69 54.45 59.23 57.28 51.13
VCD 52.38 58.04 57.86 60.85 54.26 48.94
REVERIE 40.67 52.66 41.9 53.03 37.88 41.32
Haloquest 70.57 64.89 72.73 63.00 55.20 47.45
Nullu 51.99 57.90 59.22 61.78 55.98 53.17
Octopus 52.12 57.73 53.52 58.71 46.59 47.55
Ours 78.48 68.13 77.37 61.61 61.73 60.79
Table 5: Comparison on existing methods.
Conclusion
@@
In this paper, to our best knowledge, we first reveal MLLM verb hallucination and build a benchmark to probe it from various perspectives. Our experiment reveals that MLLMs suffer from severe verb hallucination in many ways, and existing training-free hallucination mitigation methods fail. Fine-tuning is still the most promising way. However, how to fine-tune existing models efficiently is still a problem to be explored. Experiments show that MLLM verb hallucination is quite different from object hallucination. Moreover, whether there are effective training-free verb hallucination mitigation methods is still an open question.
Acknowledgements
This work was supported in part by the National Natural Science Foundation of China under Grant No. 62306175, the Shanghai Committee of Science and Technology, China (Grant No. 24511103200), the National Key Research and Development Project of China (No. 2022ZD0160102), and Shanghai Artificial Intelligence Laboratory, XPLORER PRIZE grants.
@@ Appendix A Details on Benchmark Construction
Benchmark Based on HICO
HICO (chao2015hico) test set contains 9.8K images. We use the HICO test set for benchmark construction. It includes 600 action classes formed by 80 object classes and 117 verb classes. There is a special verb class “no_interaction”, and we leave out annotations about it.
@@
For MC question construction, given an image $I$ and its HOI annotation $\mathcal{H}_{I}$, we select $h_{0}=(v_{0},o_{0})\in\mathcal{H}_{I}$ as a positive option, select 3 negative options from the set of plausible but not presented action classes $\{(v,o)\in\mathcal{H}|o=o_{0}\land(v,o)\notin\mathcal{H}_{I}\}$, and form a MC question with one correct option. In our benchmark, there are 16.6K unique MC questions.
@@
Benchmark Based on CharadesEgo
CharadesEgo (sigurdsson2018actor) contains 157 action classes, each of which is formed by a verb and an object. However, many object classes in CharadesEgo can not match more than 2 verbs. Therefore, to build negative options, we sample from the set $\{(v,o)\in\mathcal{H}|o\in\mathcal{O}_{I}\land(v,o)\notin\mathcal{H}_{I}\}$, where $\mathcal{O}_{I}=\{o|(v,o)\in\mathcal{H}_{I}\}$. The ratio of questions with correct answers Yes and No is approximately 1:1. There are 76K unique YN questions and 36K unique MC questions, and we use 28K YN questions and 10K MC questions for experiment in the main paper.
@@ Appendix B Additional Explanation and Analysis on Model Behaviors
of LLaVA V1.5 (liu2024improved). In our setting, $y_{0}$’s are usually option tokens or Yes/No tokens, which are used to judge the correctness of MLLMs’ answers. Here we provide some additional visualizations in Fig. 12. The visualizations give us more solid proof that uncertainty is strongly related to MLLM verb hallucination.
Figure 12: Token uncertainty of InstructBLIP (dai2023instructblip). Not hallucinated tokens are often given with high probability.
Key Image Area Attention
LLaVA V1.5 adopts CLIP (radford2021learning) as the visual encoder. 576 tokens generated by CLIP are used as visual tokens and fed to the language model after projection. Each of these 576 tokens represents a small area of the original image. Using the bounding box annotation of HICO, we can get the key image area of a certain verb, represented by the area bounded by the bounding boxes of humans and objects. Our visualization shows that hallucinated MLLM pays less attention to the key image area, but the margin is not large enough. Although MLLM can pay attention to the correct key areas, it can not often understand verb concepts.
Verb Token Attention
@@
Visual Token Attention
Visual Token Attention(VTA) is defined similarly to Visual Modality Contribution(VMC) (chen2024multi). The Visual Token Attention is defined as
@@
where $\alpha_{ij}$ represents the attention weight assigned to token $i$ at head $j$ in the last transformer layer, $V$ denotes the set of visual tokens, and $T$ denotes the set of textual tokens. In the implementation of OPERA (huang2024opera), when generating the first 10 tokens, a reward is given for high visual token attention. We visualize VTA on other models and other question types in Fig. 13 and  14, and have the following observations:
@@
Object Reliance
We split YN question sets according to object classes, and show the difference between mPLUG-Owl2’s (ye2024mplug) accuracy with and without object references in Fig. 15. We can see that for questions about most object classes, mPLUG-Owl2 behaves differently with and without object references and thus is affected by object references although the accuracy does not change much (62.94/62.61).
@@
Qwen-VL-Chat mPLUG-Owl2 LLaVA V1.5 InstructBLIP
overlap 95.35 79.17 82.14 100
Table 6: Influential object set overlap of MLLMs.
IoU Qwen-VL-Chat LLaVA V1.5 mPLUG-Owl2 InstructBLIP
Qwen-VL-Chat - 59.30 42.86 52.78
LLaVA V1.5 59.30 - 46.91 53.57
mPLUG-Owl2 42.86 46.91 - 41.54
InstructBLIP 52.78 53.57 41.54 -
Table 7: YN influenced verb set IOU.
MLLM Error Consistency
[…]
Consistency among models. We analyze models and get verb classes that are influenced by object references most according to per verb class Yes ratio. From Tab. 7, we can see that there is substantial overlap among different models.